OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (texaschic101):

first you need to find the slope of the equation given. We do this because the slope of the given equation is the negative reciprocal of the perpendicular line. Do you know how to find the slope of : y = -4x - 2 ??

OpenStudy (texaschic101):

In y = mx + b form, the slope is in the m position . y = mx + b y = -4x - 2.....now what number is in the m position ??

OpenStudy (anonymous):

-4 is the slope

OpenStudy (texaschic101):

good. But for a perpendicular line we need the negative reciprocal slope. All that means is " flip " the slope and change the sign. So we will flip -4 and change the sign, making the slope we need be 1/4. Do you see how I flipped the slope and changed the sign ?? Understand so far ?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

and now I do y+11 = 1/4 (x+16)

OpenStudy (anonymous):

I change the x and find the y, right ?

OpenStudy (texaschic101):

now use y = mx + b slope(m) = 1/4 (-16,-11)...x = -16 and y = -11 now we sub and find b, the y intercept -11 = 1/4(-16) + b -11 = - 4 + b -11 + 4 = b - 7 = b so our perpendicular equation is : y = 1/4x - 7 any question ?

OpenStudy (anonymous):

I have another one that is similar but i will try to so it my self. Ill ask you if I have questions, THANK YOU!

OpenStudy (texaschic101):

or this way.. y - y1 = m(x - x1) slope(m) = 1/4 (-16,-11)..x = -16 and y = -11 now sub y - (-11) = 1/4(x - (-16) y + 11 = 1/4(x + 16) y + 11 = 1/4x + 4 y = 1/4x + 4 - 11 y = 1/4x - 7 either way should get you to the right answer

OpenStudy (anonymous):

okay thank you